OpenStudy (anonymous):

you are planning to clean the sky lights on yur roof. the rof line is 9 feet off the ground and your ladder is 16feet long. the ladder needs to extend 6 feet beyond the roof line in order to let you access the roof safely. how far from the house do you need to place the ladder? round t0 th nearest foot

OpenStudy (anonymous):

16^2 = 15^2 + x^2 256-225 = x^2 x^2 = 31 x = 6 feet

OpenStudy (gw2011):

I believe the answer should be 13 feet which is rounded to the nearest foot. Here is how I arrived at the answer. Since the ladder is 16 feet long and 6 feet of this ladder has to extend beyond the roof line, then you have 10 feet of the ladder remaining for placing. This will give you a right triangle which is solved asa follows: c^2=a^2+b^2 (10)^2=a^2+(9)^2 a^2=(10)^2-(9)^2 a^2=100+81 a^2=181 a=13.45 feet or rounding to the nearest foot a=13 feet

OpenStudy (anonymous):

both are wrong. the answer is sqrt(100-81)=sqrt(19) feet

OpenStudy (anonymous):

which gives about 4 feet

OpenStudy (gw2011):

Sorry, you're right--it should be about 4 feet. I made a mistake in my sign. I originally showed (10)^2-(9)^2 but then in the next line I accidently showed 100+81 which should have been 100-81.

OpenStudy (anonymous):

yes. :) we all make mistakes